A vitamin syrup for infants contains 27.44 mg of zinc sulfate monohydrate (ZnSO4.H2O) and 100.00 mg of ascorbic acid (C6H8O6)for every 5 ml.

How many moles of asorbic acid does it contain?
1
Expert's answer
2021-10-01T09:27:04-0400

n = m / M

M (C6H8O6) = 176.12 g/mol

100 mg = 0.1 g

n (C6H8O6) = 0.1 / 176.12 = 0.0006 mol

A vitamin syrup contains 0.0006 mol of ascorbic acid C6H8O6 for every 5 ml, or 0.11 mol of ascorbic acid C6H8O6 for 1 liter.